a train 120 m long passes a railway platform 160m long in 14 sec . how long will it take to pass another platform which is 100 m long ?
 

Grade:12

1 Answers

Neeti
8 years ago
in the first case, the total distance travelled by the train is 160 + 120 = 280
time taken is 14sec therefore speed = distance / time = 280 / 14 = 20m/s
 
Now in the second case, the distance travelled by the train is 100 plus 120 = 220m and speed of the train is 20m/s
 
therefore time = distance/ speed = 220 / 20 = 11seconds.
 
Please approve if you like and understand my solution and feel free to ask any other doubt :)